OpenStudy (anonymous):

10. A coat is on sale for Rs.120 after a discount of 20 percent. Find the original price.?

OpenStudy (nirmalnema):

if there is the discount of 20 % then 120 is 80 of original price. from here u can find original price let the original price be x 120/x *100 = 80 x=12000/80 =150

OpenStudy (anonymous):

120/x *100 = 80 could u plz justify it more accurately?

OpenStudy (anonymous):

what does the sign placed above 100 suggest?

OpenStudy (nirmalnema):

multiplication

OpenStudy (anonymous):

thanks

OpenStudy (nirmalnema):

welcome

OpenStudy (anonymous):

(1-0.20) C = 120 solve for C, the original price of the coat